Given the supply function and inverse demand function, Qs = 40 + 20 P and p=10-1/10Q

respectively. Determine the equilibrium price and quantity?


1
Expert's answer
2021-04-29T10:28:10-0400
"Q_S=40+20p"

"Q_D=100-10p"

"100-10p=40+20p"

"30p=60"

"p=2"

"Q=80"